# Tying knots in particle physics

### v1

@@ Introduction
Now, make the $\phi_{1}$ and $\phi_{2}$ strings linked. Particularly, one simple example is a link made of a single $\phi_{2}$ string and multiple $\phi_{1}$ strings (See the left panel of Fig. 2). Now the coupling (3) plays a crucial role; it induces the electric flux of the $A_{\mu}$ field around the $\phi_{1}$ string. To see this, let us rewrite the Chern-Simons coupling as
$\mathcal{L}_{\mathrm{CS}}$ $=-CA^{0}\,\left(\vec{\nabla}a\cdot\vec{B}\right)\,,$ (4)
@@
Now we apply our study on the knot soliton to particle physics models and discuss its influence on cosmology. One realistic and promising setup is obtained by identifying the $U(1)_{\mathrm{local}}$ and $U(1)_{\mathrm{global}}$ symmetries in (1) as the $U(1)_{B-L}$ and $U(1)_{\mathrm{PQ}}$ symmetries, respectively. In order for $\phi_{1}$ to couple with right-handed neutrinos, the $B-L$ charge of $\phi_{1}$ is taken as $q_{1}=2$ instead of $1$ while keeping $q_{2}=0$. We emphasize that this setup is well motivated because it can elegantly resolve some of the mysteries in particle physics as stated above.101010 When one allows $\phi_{2}$ to have a small non-zero charge under $U(1)_{B-L}$ and introduces fermionic matters charged under the $U(1)_{B-L}$ and $U(1)_{\mathrm{PQ}}$ symmetries, the setup may have a bonus Fukuda:2017ylt ; Ibe:2018hir that the quality of the $U(1)_{\mathrm{PQ}}$ symmetry is protected against the effect of quantum gravity. (See Refs. Fukuda:2017ylt ; Ibe:2018hir for more details.) In this setup, the QCD axion is identified with the phase of $\phi_{2}$. For this to solve the strong CP problem, we need additional fermionic fields inducing the Chern-Simons coupling between the QCD axion and gluon, which may be realized by introducing Kim-Shifman-Vainshtein-Zakharov(KSVZ)-like Kim:1979if ; Shifman:1979if heavy quarks. How the QCD axion couples to the $B-L$ and SM gauge bosons depends on the charge assignment of the fermionic sectors, which is kept as general as possible hereafter. From the astrophysical constraints on the QCD axion, we take $v\gtrsim 10^{8}\,\mathrm{GeV}$ Raffelt:2006cw ; Irastorza:2018dyq ; Chang:2018rso ; Hamaguchi:2018oqw .
@@
Now let us discuss cosmological fate of the knot solitons. The Universe was very hot and full of thermal relativistic particles at early stages (after the cosmic inflation). When the temperature $T$ of the Universe is much larger than $v$, the $U(1)_{B-L}$ and $U(1)_{\mathrm{PQ}}$ symmetries are not broken while they get spontaneously broken almost simultaneously due to a phase transition when $T$ gets less than a critical temperature $T_{c}\simeq v$. At the transition, the $\phi_{1}$ and $\phi_{2}$ strings are produced by the Kibble-Zurek mechanism Kibble:1980mv ; Zurek:1985qw . The produced strings have completely random configurations, in which some of the strings are linked with finite probability. After the oscillation and the random motion are relaxed, long $\phi_{1}$ and $\phi_{2}$ strings form conventional networks whose typical length scale is the Hubble scale. Besides them, small loops with the linking configurations become the knot solitons that we have studied above. In Ref. Vachaspati:1994ng , the linking number density of a single species of string formed during a phase transition is given as $10^{-4}\xi^{-3}$, where $\xi$ is the correlation length to be $\xi\sim T_{c}^{-1}$. Thus we use this value as a reference value for the number density of the (anti-)knot solitons: $n_{\mathrm{knot}}(T=T_{c})\sim 10^{-4}\xi^{-3}$. [^11] Note that results given below do not depend on the value of the produced number density $n_{\mathrm{knot}}(T_{c})$ as long as they dominate the energy density of the Universe before they decay.
@@
Our Universe is baryon asymmetric, i.e., baryon is more than anti-baryon as characterized by a quantity $Y_{B}\equiv(n_{B}-n_{\bar{B}})/s\simeq 0.8\times 10^{-10}$ Planck:2018vyg ; ParticleDataGroup:2022pth where $n_{B(\bar{B})}$ and $s$ are the number density of (anti)baryon and the entropy density, respectively. The SM however cannot explain this asymmetry. As stated above, the knot solitons remain in the Universe and later decay into particles. When $T_{rh}$ is larger than the electroweak (EW) scale $\mathcal{O}(10^{2})$ GeV, 131313 There can be an alternative scenario of baryogenesis even when $T_{rh}$ is smaller than the electroweak scale. This is complementary to the scenario presented here. See Appendix .3 for the details. this decay process can produce baryon asymmetry in the Universe in a similar way to non-thermal leptogenesis via inflaton decay Asaka:1999yd ; Giudice:1999fb ; Asaka:1999jb ; Lazarides:1990huy ; Kumekawa:1994gx . Note that the baryon asymmetry produced at earlier stages than the end of the knot domination is diluted by the secondary reheating due to the knot decay. For simplicity, we here assume it to be sufficiently diluted away and concentrate on the baryon asymmetry produced by the knot decay.
@@ Acknowledgments
To get the observed baryon asymmetry, we need the mass of the knot soliton $M$ smaller than $10^{12}$ GeV. However, $M$ is typically given as $M\sim\mathcal{O}(10^{3})\,v/g$, and $v$ should satisfy $v\gtrsim 10^{8}$ GeV, in which the last inequality comes from astrophysical constraints on the axion decay constant, $f_{a}\gtrsim\mathcal{O}(10^{8})$ GeV Raffelt:2006cw ; Irastorza:2018dyq ; Chang:2018rso ; Hamaguchi:2018oqw . Thus successful baryogenesis works for $10^{8}\,\mathrm{GeV}\lesssim v\lesssim 10^{9}\,\mathrm{GeV}$.

### v2

@@ Introduction
Now, make the $\phi_{1}$ and $\phi_{2}$ strings linked. Particularly, one simple example is a link made of a single $\phi_{2}$ string and multiple $\phi_{1}$ strings (See the left panel of Fig. 2). Now the coupling (3) plays a crucial role; it induces the electric flux of the $A_{\mu}$ field around the $\phi_{1}$ string. To see this, let us take a functional derivative of the Chern-Simons coupling with respect to $A_{0}$ as
$\frac{\delta}{\delta A_{0}}\int d^{4}x\,\mathcal{L}_{\mathrm{CS}}$ $=-C\,\left(\vec{\nabla}a\cdot\vec{B}\right)\,,$ (4)
@@
Now we apply our study on the knot soliton to particle physics models and discuss its influence on cosmology. One realistic and promising setup is obtained by identifying the $U(1)_{\mathrm{local}}$ and $U(1)_{\mathrm{global}}$ symmetries in (1) as the $U(1)_{B-L}$ and $U(1)_{\mathrm{PQ}}$ symmetries, respectively. In order for $\phi_{1}$ to couple with right-handed neutrinos, the $B-L$ charge of $\phi_{1}$ is taken as $q_{1}=2$ instead of $1$ while keeping $q_{2}=0$. We emphasize that this setup is well motivated because it can elegantly resolve some of the mysteries in particle physics as stated above.101010 When one allows $\phi_{2}$ to have a small non-zero charge under $U(1)_{B-L}$ and introduces fermionic matters charged under the $U(1)_{B-L}$ and $U(1)_{\mathrm{PQ}}$ symmetries, the setup may have a bonus Fukuda:2017ylt ; Ibe:2018hir that the quality of the $U(1)_{\mathrm{PQ}}$ symmetry is protected against the effect of quantum gravity. (See Refs. Fukuda:2017ylt ; Ibe:2018hir for more details.) In this setup, the QCD axion is identified with the phase of $\phi_{2}$. For this to solve the strong CP problem, we need additional fermionic fields inducing the Chern-Simons coupling between the QCD axion and gluon, which may be realized by introducing Kim-Shifman-Vainshtein-Zakharov(KSVZ)-like Kim:1979if ; Shifman:1979if heavy quarks. How the QCD axion couples to the $B-L$ and SM gauge bosons depends on the charge assignment of the fermionic sectors, which is kept as general as possible hereafter. From the astrophysical constraints on the QCD axion, we take $v\gtrsim 10^{8}\,\mathrm{GeV}$ Irastorza:2018dyq ; Chang:2018rso ; Hamaguchi:2018oqw ; Caputo:2024oqc .
@@
Now let us discuss cosmological fate of the knot solitons. The Universe was very hot and full of thermal relativistic particles at early stages (after the cosmic inflation). When the temperature $T$ of the Universe is much larger than $v$, the $U(1)_{B-L}$ and $U(1)_{\mathrm{PQ}}$ symmetries are not broken while they get spontaneously broken almost simultaneously due to a phase transition when $T$ gets less than a critical temperature $T_{c}\simeq v$. At the transition, the $\phi_{1}$ and $\phi_{2}$ strings are produced by the Kibble-Zurek mechanism Kibble:1980mv ; Zurek:1985qw . The produced strings have random configurations, in which some of the strings are linked with finite probability. After the oscillation and the random motion are relaxed, long $\phi_{1}$ and $\phi_{2}$ strings form conventional networks whose typical length scale is the Hubble scale. Besides them, small loops with the linking configurations become the knot solitons that we have studied above. Let us estimate the number density of the knot solitons produced by this mechanism. One should note that, when $\kappa$ is sufficiently small compared to the other parameters $\lambda,g$ and $C$, the scalar potential $V$ has an approximate global $O(4)$ symmetry spontaneously broken down to $O(3)$. Thus, within a microscopic length scale smaller than $1/(\sqrt{\kappa}T_{c})$, the linking number density is well approximated by the number density of textures Turok:1989ai ; Borrill:1991mv ; Leese:1991gt ; Vilenkin:2000jqa rather than that of linking cosmic strings Vachaspati:1994ng . According to Ref. Turok:1989ai , the linking number density is given as $\sim 0.04\,\xi^{-3}$, where $\xi$ is the correlation length to be $\xi\sim T_{c}^{-1}$. Noting that the knot solitons that we found above should have the linking number larger than or equal to $4$, the number density of the (anti-)knot solitons is estimated as $n_{\mathrm{knot}}(T=T_{c})\sim(0.04)^{4}\,\xi^{-3}$. [^11] Note that results given below do not depend on the value of the produced number density $n_{\mathrm{knot}}(T_{c})$ as long as they dominate the energy density of the Universe before they decay.
@@
Our Universe is baryon asymmetric, i.e., baryon is more than anti-baryon as characterized by a quantity $Y_{B}\equiv(n_{B}-n_{\bar{B}})/s\simeq 0.8\times 10^{-10}$ Planck:2018vyg where $n_{B(\bar{B})}$ and $s$ are the number density of (anti)baryon and the entropy density, respectively. The SM however cannot explain this asymmetry. As stated above, the knot solitons remain in the Universe and later decay into particles. When $T_{rh}$ is larger than the electroweak (EW) scale $\mathcal{O}(10^{2})$ GeV, 131313 There can be an alternative scenario of baryogenesis even when $T_{rh}$ is smaller than the electroweak scale. This is complementary to the scenario presented here. See Appendix .3 for the details. this decay process can produce baryon asymmetry in the Universe in a similar way to non-thermal leptogenesis via inflaton decay Asaka:1999yd ; Giudice:1999fb ; Asaka:1999jb ; Lazarides:1990huy ; Kumekawa:1994gx . Note that the baryon asymmetry produced at earlier stages than the end of the knot domination is diluted by the secondary reheating due to the knot decay. For simplicity, we here assume it to be sufficiently diluted away and concentrate on the baryon asymmetry produced by the knot decay.
@@ Acknowledgments
To get the observed baryon asymmetry, we need the mass of the knot soliton $M$ smaller than $10^{12}$ GeV. However, $M$ is typically given as $M\sim\mathcal{O}(10^{3})\,v/g$, and $v$ should satisfy $v\gtrsim 10^{8}$ GeV, in which the last inequality comes from astrophysical constraints on the axion decay constant, $f_{a}\gtrsim\mathcal{O}(10^{8})$ GeV Irastorza:2018dyq ; Chang:2018rso ; Hamaguchi:2018oqw ; Caputo:2024oqc . Thus successful baryogenesis works for $10^{8}\,\mathrm{GeV}\lesssim v\lesssim 10^{9}\,\mathrm{GeV}$.